Proposal to the MIT Alumni Office for Alumni Dialin/Login Service

BLEED ON ME!!!

Derek A. Atkins Chairman, Student Information Processing Board

1 The Problem

Currently, when a student graduates from MIT, unless he or she becomes an MIT graduate student or acquires another form of affiliation with MIT, all contact with the Institute is lost, except through the Alumni Office. One of the services lost upon exiting from the Institute is access to services like those provided by Athena, namely access to computer services. Is there a way for MIT to provide a service to its alumni similar to that which Athena provides to the current students?

Athena is MIT’s computational environment, and every student is allowed to get their own Athena account, which lets them communicate with students and others both at MIT and outside of MIT. Access to Athena gives students a place to work: a place to work on papers, to talk to friends, to collaborate on work with a professor, to communicate to the educational and professional world at large.

There are a number of alumni each year who, for one reason or another, would like to keep their access to Athena. The current Athena policy does not allow for this at all, since only students can have an Athena accounts, so the Alumni are forced to try to find a new method to obtain those services. Providing such a service would allow Alumni, for example, to keep communicating with each other after they leave, or to keep in contact with those organizations with whom they were involved when they were at MIT.

This document is a proposal to the Alumni Association to help fund the service described herein: an Alumni Server to allow electronic access to alumni.

2 The Solution

What we propose is to create a new service available to MIT that would allow alumni access to online electronic services similar to what Athena provides to the MIT community. We hope that this is not limited in scope to only those Alumni in the Boston area, although given the resources required to expand to other areas it seems feasible to begin locally.

An example of the services that we propose to support include:

E-mail: The ability to send and receive electronic mail.

NetNews: The ability to read and post messages to the Usenet.

Network Access: The ability to access any host on the Internet, using:

– telnet or rlogin: To access accounts at other sites or on other machines.

– ftp, the File Transfer Protocol: To transfer files.

File Transfer: The ability to download personal mail/files and freely distributible software to computers at one’s home or office.

Disk Space: User storage space for personal programs or files. Each user would be allotted a limited amount of space when they subscribe to the service.

Zephyr: The ability to communicate in real time to other users of the system. This might be tied into the Athena Zephyr system, depending on feasibility.

Other: The ability to access all Athena services which are already available at remote sites, Discuss and AFS, for example.

3 The Requirements

The major requirements for this project are machines and disk space. We would need at a minimum two machines: one to act as a server for all the standard Athena network services, and the other to act as a computron/login server. In addition, a 4mm (DAT) tape drive would be useful in ensuring reliable backups. The amount of disk space, memory, and the number of machines necessary are dependent upon the load that we anticipate from usage.

As an example, Athena is currently using approximately 8 dialup machines. Three of them are IBM PC/RT machines and the other five are VAXstation 3100s. According to Athena:

As to the hardware; there are 5 VS3100’s and 3 RT’s. The Vaxen have 20M of memory, and total of 280M of disk (rz55’s), using 80M for swap. The RT’s have 12M of memory, using Athens drives, and have 57M for swap.

These are just the login server configurations, and do not take into account the space used up for file storage and maintenance databases.

As for usage, looking at Table 1 shows a week’s worth of usage on the Athena dialup machines. Dividing through by the eight machines, it can be seen that the average usage is about 10-15 users per machine, continuously, although the loading of the machines will vary depending on the time of year.

Athena is supporting a user base of approximately 5000-10000 users. We propose to start a lot smaller than that, estimating less than 1000 users at the start. Theoretically this means we could get by with only two moderately loaded machines.

We doubt that we would need the hardware support that Athena does. Using two or three machines for filespace and another five or six for login ability would more than equal the Athena login ability, from the perspective of computer resources per user.

A good medium solution would be to start with two fileserver machines and three dialup machines. This would allow us a starting base well within our goals, and allow us to expand the system for a while without additional need, until the usage base grows, requiring more hardware.

As for other requirements, we would also need to deal with incoming phone lines and a CPT (Cisco Protocol Transformer (sp?)) so that we do not abuse Athena’s resources for dial-in capability.

The Alumni Office would sponsor this project, and would pay for the hardware and physical maintenance costs. The SIPB would provide for the administrative needs of the system for two years, at which point a decision should be made; considering the load and usage whether or not the SIPB

                Load            Simult. Users
Date      Min    Max    Avg.    Min Max Avg.         % up
---------------------------------------------------------
05/10      0.00  20.58 1.0129    14 126 78.33        87.5
05/11      0.02   9.88 1.2014    38 206 116.8     92.7083
05/12      0.00  11.07 1.3926    34 204 120.6         100
05/13      0.00  13.90 1.4308    34 211 111.6       93.75
05/14      0.01  11.77 1.3771    25 194 110.9         100
05/15      0.00   5.81 0.8907    31 191   100         100
05/16      0.00   5.07 0.6401    25 104 63.12         100

                     Load            Simult. Users
Machine        Min    Max    Avg.    Min Max Avg.         % up
--------------------------------------------------------------
e40-008-10      0.00   8.12 1.0456     2  31 13.66         100
e40-008-11      0.00   5.44 0.9045     1  34 12.85         100
e40-008-13      0.01  20.58 1.4936     1  34 12.15         100
e40-008-5       0.00  11.07 1.1068     4  37 15.52         100
e40-008-6       0.01   9.53 1.1098     1  29 9.005     77.3809
e40-008-7       0.00   8.82 1.0857     2  31 13.07         100
e40-008-8       0.00  11.77 1.2963     0  37 11.63     92.8571
e40-008-9       0.00  13.90 1.0384     0  36 12.29         100

Table 1: Athena Dialup Usage Statistics for the week ending Sunday, May 17, 1992 could actually continue running the service and still provide adequate quality of service.

4 The Benefits

The service would be Athena-like and would provide a majority of the services of Athena, and, we hope, would be able to provide some of Athena’s thirdparty software as well.

One place look for a precedent is at existing companies which supply dialup network access similar to what we are proposing to supply. They do not supply many of the MIT-specific services we intend to supply, however. For example, World, which is run by Software Tool and Die here in the Boston area, charges approximately $20.00 per month, and offers access to the Internet at large. Well, on the other hand, which is located in San Francisco, does not offer all the services of World. Cost? The Alumni Office could easily decide to charge users for their accounts and also for extra storage space.

Another service that could be provided is an online directory of alumni, which would make it easier for both the alumni and the Alumni Office to keep track of alumni. To prevent improper use of this information, the number of accesses per user per time could be limited or access to the database could be prohibited to external sites. This could be used as an online white pages of MIT Alumni. We know that the Alumni Office has spoken with IS/Athena regarding this service, however we think that conglomerating many alumni services into our proposed system would be offer the best possible service to its users.

5 Possible Problems

Some problems that we might have include:

Loading If we were oversubscribed then we would need to expand our base of machines. We do not anticipate this happening at the beginning of the experiment.

Software Licenses Who will pay for third-party software? Many MIT site licenses are for educational use only. How can we be sure that alumni will use the software for educational purposes? Otherwise we might have to buy licenses ourselves, which could get expensive. Should we just ignore this issue? Who will pay for it? The growing availability of good free software may nullify this problem shortly. Since this venture will be funded through the Alumni Association, the licenses should be bought through the Alumni Office. If the licenses are for educational or academic use only, then we will have to limit the availability of the software, or try to work out an arrangement with the software vendor in order to provide those services.

Overloading Services If we started out by using Athena’s servers, we might wind up loading them even more than they are now which means we would have to purchase, set up, and maintain our own servers to handle our own load. One such service that falls under this category is the Athena Zephyr notification service. It is currently loaded almost to capacity from the normal Athena usage, and adding more load through the Alumni Server might overload this service. This hopefully would not be a problem in the beginning, but it might become a problem if usage increased, which means we would have to set up our own server, which would take some resources.

NetNews Licenses Some of the newsgroups Athena receives are at a cost, and there are license agreements, as with ClariNet, which prohibit reproduction. This would have to be worked out for such newsgroups to be made available to our users. We have to decide a policy regarding these groups, whether we want to purchase them in this service, or bind the users to the agreements when they use the service.

There will undoubtedly be numerous other problems with this system that we have not foreseen, but the fact that this has not been undertaken in the past makes it an exciting a worthwhile task to undertake. We hope that people would become interested in it as word of its existence spreads among the alumni. In casual contact with graduating seniors and other new alumni, we have asked their opinions regarding this proposal, and the straw poll seems to be positive. Many recent alumni, it appears, don’t want their accounts to disappear, or at least want some sort of access if it does. This service will provide the alumni with that opportunity.